OpenStudy (anonymous):

Of the new cars in a car dealer's lot, 1 in 6 are white. Today, 4 cars were sold. A) What is the probability that 3 of the cars sold were white? B) What is the probability that at least 2 of the cars sold were white?

OpenStudy (anonymous):

@tcarroll010 Do you know anything about probability? lol

jimthompson5910 (jim_thompson5910):

Are you familiar with the binomial probability theorem?

OpenStudy (anonymous):

Not really

jimthompson5910 (jim_thompson5910):

so you haven't learned it yet in class?

OpenStudy (anonymous):

No, he hasn't taught us that.

OpenStudy (anonymous):

Wait, I think I know what you're talking about.

jimthompson5910 (jim_thompson5910):

oh you have heard about it?

OpenStudy (anonymous):

I think I've seen it before, but we have never used it in class.

jimthompson5910 (jim_thompson5910):

so you've seen in your book then?

OpenStudy (anonymous):

Yeah

jimthompson5910 (jim_thompson5910):

So you've seen the following formula before \[\Large P(X = k) = _{n}C_{k}p^k(1-p)^{n-k}\]

OpenStudy (anonymous):

I've seen it, but I'm not quite sure how it works.

jimthompson5910 (jim_thompson5910):

In this case, p = 1/6 since the probability of selecting a white care on each trial is 1/6

jimthompson5910 (jim_thompson5910):

n = 4 because there are 4 cars sold (ie 4 trials done)

jimthompson5910 (jim_thompson5910):

and for part a) k = 3 because we want to know the probability of selling exactly k = 3 white cars

jimthompson5910 (jim_thompson5910):

So \[\Large P(X = k) = _{n}C_{k}p^k(1-p)^{n-k}\] becomes \[\Large P(X = 3) = _{4}C_{3}\left(\frac{1}{6}\right)^{3}\left(1-\frac{1}{6}\right)^{4-3}\]

OpenStudy (anonymous):

I'm writing this down as we go. I got all of that so far. What do you do next?

jimthompson5910 (jim_thompson5910):

Then, you use your calculator (or you can compute by hand, let me know if you want to see it by hand) to compute 4 C 3 4 C 3 = 4 Then you use a calculator to compute 4*(1/6)^(3) * (1 - 1/6)^(4-3) to get 0.01543209876543 So \[\Large P(X = 3) \approx 0.01543209876543\]

OpenStudy (anonymous):

Side note: Jim I know you know what I'm about to say, so this if for Haley's benefit. In the use of the binomial theorem, independence is assumed, meaning that each event is independent of the outcome of the other events. Aside from the math, so you can understand the application of this formula, there is a big assumption going on and that is that the dealer has a HUGE number of cars so that 1/6 and 5/6 are approximated as close to that for each trial repetition. Strictly speaking, the binomial theorem can be used here with the understanding that it becomes an approximation only because of the huge number of cars. If the dealer had, say 40 cars, the binomial would be too far off. If the dealer has like 500 cars, the binomial gives a good approximation. This is a problem in "without replacement", so caution is to be taken with results in that approx is assumed. I'm sure your teacher wants to use this here because we are not given the # of cars, just assuming a LOT.

OpenStudy (anonymous):

I think I got it now! Thanks! He never taught us this before, I'm not sure why though.

OpenStudy (anonymous):

Do you use the same formula for the second problem?

jimthompson5910 (jim_thompson5910):

yes, but you need to compute P(X = 2) P(X = 3) P(X = 4) then add up the 3 individual probabilities

jimthompson5910 (jim_thompson5910):

n and p will remain the same

OpenStudy (anonymous):

Is k=2?

jimthompson5910 (jim_thompson5910):

for P(X = 2), k = 2, yes

jimthompson5910 (jim_thompson5910):

for P(X = 3), k = 3, etc etc

OpenStudy (anonymous):

I tried it and I messed up somewhere. I'm still kind of confused with the formula.

jimthompson5910 (jim_thompson5910):

what did you get for P(X = 2)

OpenStudy (anonymous):

I'm pretty sure its wrong, but I got .0231481481

jimthompson5910 (jim_thompson5910):

you should have gotten 0.1157407

OpenStudy (anonymous):

I'm going to try again. N= 4 P=1/6 K=2, right?

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

I got 0.1157407407 for P(X=2) now. I'm going to redo P(X=4)

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

So far I have P(X=4)=1*1/6^4*(1-1/6)^(4-4) Is that right?

jimthompson5910 (jim_thompson5910):

so far, so good

OpenStudy (anonymous):

For 1/6^4 I got this....

jimthompson5910 (jim_thompson5910):

that can be written as 0.0007716

jimthompson5910 (jim_thompson5910):

either are correct

jimthompson5910 (jim_thompson5910):

now add up what you got for P(X = 2), P(X = 3), and P(X = 4)

OpenStudy (anonymous):

I got .13

jimthompson5910 (jim_thompson5910):

or, more accurately, 0.131944444444444 but if you round to the nearest hundredth, then you'll get 0.13

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

yw